OpenStudy (anonymous):

x^​‑3*​x^​7*​x Simplify the expresson using each base only once

OpenStudy (raden):

hint : x^a * x^b * x^c = x^(a+b+c)

OpenStudy (anonymous):

i dont understand what ur tryingg to say ....

OpenStudy (raden):

just adding the powers, because the bases already be same

OpenStudy (raden):

well, this is an example for you x^-2 * x^8 * x = x^(-2 + 8 + 1) = x^7 so, x^7 its answer

OpenStudy (anonymous):

thanks .... \[1/r^-3 \] is 1r-3 right?

OpenStudy (anonymous):

^^^ where did u get 8 from the numbers are -3 7

OpenStudy (anonymous):

and 1

OpenStudy (anonymous):

i got the awnser for the x though so never mind bout the 8 .... um can you help me with a nother proble ???

OpenStudy (raden):

x^-2 -----> -2 as the power/exponent x^8 -----> 8 as the power x ------> the power is 1 x is the base

OpenStudy (raden):

use this property : 1/a^-m = a^m

OpenStudy (anonymous):

there wasnt no 8 thats what i was trying to tell u.

OpenStudy (anonymous):

(t^-3)^3 how i do this?? multiply the outer 3 by negitive 3 ? t^-9 >>>change it 1 / t^9 ????

OpenStudy (raden):

yes, that's right!

OpenStudy (anonymous):

Thank :)))

OpenStudy (raden):

welcome :)